# Revisiting Few-Shot Object Detection with Vision-Language Models

### v2

@@ 3.3 FSOD Benchmarks are Federated Datasets
Table 1: VLM Zero-Shot Inference Is a Strong FSOD Baseline. Zero-shot inference with VLMs like GroundingDINO resoundingly outperform state-of-the-art FSOD methods on the COCO FSOD benchmark, motivating the need to re-frame FSOD to embrace foundation models.
@@
AP bAP nAP
FRCN-ft-full [46] 18.6 20.6 12.5
FRCN-BCE [46] 30.2 36.8 10.3
TFA w/ fc [38] 29.3 34.5 13.5
TFA w/cos [38] 29.9 35.3 13.6
MPSR [42] 17.1 18.1 14.1
Meta-RCNN [46] 7.8 7.1 9.1
FsDetView [43] 10.0 9.3 12.0
Retentive R-CNN [7] 32.9 39.3 13.8
DiGeo [25] 33.1 39.4 14.2
GroundingDINO (Zero-Shot) [23] 48.3 46.3 54.3
4 Experiments
We conduct extensive experiments to validate that zero-shot inference from VLMs significantly improves over state-of-the-art FSOD approaches, suggesting that existing benchmarks should be re-framed to include foundation models. Moreover, we demonstrate that using federated losses consistently improve fine-tuning performance on LVIS and nuImages. Lastly, we compare our performance against an (oracle) upper bound and ablate the impact of freezing different detector components on fine-tuning performance.
Datasets and Metrics. We repurpose two established datasets for Foundational FSOD:
  • LVIS [11] re-annotates COCO images using 1,230 fine-grained classes, which are divided into frequent, common and rare based on the cardinality of each class. Frequent and common classes are combined to form LVIS-base and is used for pre-training. Rare classes are used for LVIS-novel. Following [38, 25], we benchmark with LVIS v0.5 on publicly released data splits and report performance averaged across 3 splits for frequent, common, and rare groups ($AP_{f},AP_{c},AP_{r}$) on the LVIS val-set.
  • nuImages [1] annotates 18 classes, which are divided into groups with many, medium, and few examples [29, 26] and report AP for each cohort. Although this dataset is not traditionally used for FSOD, nuImages’ open-world categories like debris and pushable-pullable make it particularly challenging (even for VLMs), and is a realistic benchmark for Foundational FSOD. Unlike LVIS, nuImages is fully annotated, so we construct our Foundational FSOD benchmark using the procedure described in subsection 3.3. Since the specific $K$ examples can have a significant impact on the overall performance, we run each experiment over 10 random data splits and report the average.
@@ 4.1 Zero-Shot Inference Is A Strong FSOD Baseline
Table 2: LVIS Foundational FSOD Performance. Detic pre-trained only on LVIS-base outperforms specialized methods like TFA and DiGeo by $\sim$6 AP, without fine-tuning on rare classes. Importantly these performance improvements can be attributed to Detic’s CLIP-based classifier and demonstrates that concept leakage through language can be advantageous in data constrained settings. By leveraging our insight that FSOD benchmarks are actually federated datasets, we show that fine-tuning with pseudo-negatives improves over standard fine-tuning by 1.6 AP (30.0 vs. 31.6, ResNet-50 backbone) and $4.3AP_{r}$ (15.5 vs. 19.8, ResNet-50 backbone). Lastly, we note that simply swapping the ResNet-50 backbone to a Swin transformer significantly improves performance (without modifying training data). Notably, we find that Detic’s rare class performance improves over naive fine-tuning by $5.9AP_{r}$ (from 26.7 AP to 32.6 AP). All the methods in this table are pre-trained on LVIS-base.
Approach 10-shots
$AP$ $AP_{f}$ $AP_{c}$ $AP_{r}$
ResNet-50 Backbone
TFA w/ fc [38] 24.1 27.9 23.9 14.9
TFA w/ cos [38] 24.4 27.7 24.3 16.9
DiGeo [25] 24.9 28.5 24.6 17.3
Detic (Base Only) [49] 30.0 34.4 30.8 16.3
   + Fine-Tuning (Base + Novel) 30.0 33.2 31.9 15.5
   w/ FedLoss 30.8 33.9 32.7 17.4
   w/ InvFedloss 31.1 34.3 32.5 18.7
   w/ Pseudo-Negatives 31.6 34.8 32.8 19.8
Swin Backbone
GroundingDINO (Zero-Shot) [23]  [^2] GroundingDINO is pre-trained on Objects365, GoldG and Cap4M 27.4 32.7 23.3 18.1
Detic (Base Only) [49] 35.2 38.7 36.8 21.4
   + Fine-Tuning (Base + Novel) 35.9 37.1 37.8 26.7
   w/ FedLoss 36.5 36.7 38.3 30.4
   w/ InvFedloss 37.1 37.8 38.5 31.1
   w/ Pseudo-Negatives 37.2 37.7 38.2 32.6
@@
We compare state-of-the-art FSOD methods with zero-shot inference from GroundingDINO [23] on COCO in Table 1. Surprisingly, GroundingDINO outperforms DiGeo [25] by 16.2% AP averaged across both base and novel categories despite never being trained on COCO images. GroundingDINO’s impressive performance is due to its large-scale multi-modal pre-training on Objects365 [34], GoldG [14] and Cap4M [20]. It is worth noting that GroundingDINO achieves higher AP on novel classes than base, suggesting that novel classes in existing benchmarks (e.g. car and person) are actually not rare in the real world. Therefore, FSOD benchmarks should be re-framed to reflect real-world applications.
4.2 Foundational FSOD with LVIS
In this section, we evaluate GroundingDINO [23] and Detic’s  [49] performance on the LVIS Foundational FSOD benchmark. We train Detic from scratch on LVIS-base with a ResNet-50 backbone for fair comparison with prior work [38, 25].
As shown in Table 2, Detic outperforms all recent FSOD baselines including DiGeo [25] by about $\sim$6 points on $AP_{c}$ and $AP_{f}$ and achieves $16.3$ $AP_{r}$ without ever seeing any rare class data (e.g by prompting Detic (Base Only) with the rare class names). Importantly, these performance improvements can be attributed to Detic’s CLIP-based classifier, which has been trained with significantly more than $K$ examples of each class. This highlights the role of language in data-constrained settings.
Further, fine-tuning Detic with pseudo-negatives improves overall performance by 1.6% (30.0 vs 31.6) over naive fine-tuning. Finally, we note that simply replacing the ResNet-50 backbone with a Swin-B transformer yields a sizeable 12.8 AP improvement for rare classes (19.8 vs. 32.6).
Figure 4: Multi-Modal Fine-Tuning Improves Concept Alignment. While VLMs demonstrate strong zero-shot capabilities, they struggle when the target class is different from concepts encountered in web-scale training. Specifically, we find that GroundingDINO [23] (left) and Detic [49] (center) struggle to detect open-world categories like pushable-pullable. Fine-tuning Detic (right) with federated losses improves VLM concept alignment. Note that red boxes represent ground-truth annotations and green boxes are predictions by their respective models.
@@ 4.2 Foundational FSOD with LVIS
Table 3: nuImages Foundational FSOD Performance. We repurpose nuImages for FSOD following the dataset creation process established by [38]. We group categories by frequency into cohorts with many, medium and few examples per class (according to the fully-annotated dataset) [29, 26]. We fine-tune Detic (pre-trained on LVIS, COCO and ImageNet-21K) on $K$ examples for each of the 18 nuImages classes. We observe that existing methods like TFA perform poorly on nuImages, achieving $<3$AP. However, by simply pre-training TFA on more data and leveraging language cues via CLIP classifier, we dramatically improve performance (from 2.0 AP to 14.8 AP, 5-shots). Interestingly, accuracy across cardinalities decreases despite all classes being trained with $K$ examples. This suggests that despite pre-training on web-scale datasets, VLMs still struggle to detect rare categories like strollers, pushable-pullable, and debris, highlighting the challenge of working with nuImages.
@@
GroundingDINO (Zero-Shot) [23] 11.44 17.42 16.13 3.38
Detic (Zero-Shot) [49] 14.26 27.28 16.88 2.36
@@
   TFA [38] w/ COCO-base 1.33 2.78 1.43 0.23
   TFA [38] w/ LVIS-base 2.02 1.69 4.08 0.58
   TFA [38] w/ LVIS,IN-21K, 14.77 25.16 18.65 3.63
@@
   TFA [38] w/ COCO-base 1.21 2.55 1.19 0.31
   TFA [38] w/ LVIS-base 2.27 2.05 4.51 0.58
   TFA [38] w/ LVIS,IN-21K, 15.53 26.01 19.93 3.88
@@
   TFA [38] w/ COCO-base 1.14 2.81 0.84 0.23
   TFA [38] w/ LVIS-base 2.23 1.48 4.98 0.45
   TFA [38] w/ LVIS,IN-21K, 16.83 27.90 21.59 4.45
@@ 4.3 Foundational FSOD with nuImages
In the context of foundational models, we argue that partitioning datasets into base and novel classes no longer makes sense. Instead, FSOD methods should only train on $K$-shot annotations for $C$ target classes, and also evaluate performance on these $C$ classes. We fine-tune Detic (pre-trained on LVIS, COCO and ImageNet-21K) on $K$ examples per class and highlight model performance in Table 3. As expected, detection accuracy improves as we add more training examples. Despite large-scale pre-training, we see low accuracy for classes with few examples, highlighting the difficulty of the nuImages dataset. High intra-class variance for categories such as debris makes it difficult to generalize given few examples. According to nuImages’ annotation instructions, debris can include anything that is too big to be safely driven over, including fallen tree branches and trash bags. Similarly pushable-pullable includes trash cans, luggage, dollies, wheel barrows, and shopping carts.
@@
To contextualize our results, we evaluate TFA [38] on the nuImages Foundational FSOD benchmark. We train two variants of TFA trained on COCO-base and LVIS-base and fine-tune both models on $K$ examples of the nuImages classes. Surprisingly, both variants of TFA achieve less than $3$ AP (cf. Table 3). We posit that this is largely due to poor classification performance. Since both LVIS and COCO classes do not significantly overlap with nuImages classes, learning a classifier from few examples is extremely difficult. However, we find that simply re-training TFA with a frozen CLIP-based classifier (similar to Detic) dramatically increases performance, reiterating the utility of language and web-scale pre-training in data-constrained settings.
4.4 Oracle Performance Analysis
We compute the upper bound performance when given access to ground-truth negatives and exhaustive annotations for the few-shot data split. Recall, nuImages is exhaustively annotated, but is repurposed for Foundational FSOD.
To compute the set of ground-truth negatives for each image, we use exhaustive ground-truth annotations to determine which categories are not present. Note that this information doesn’t exist in LVIS because its ground-truth is sparsely annotated. Training with ground-truth negatives provides an upper bound on our pseudo-negatives experiment. Next, we train using exhaustive ground-truth annotations to provide an upper bound for the specific set of images used during training. In addition, this experiment highlights the performance gap between having exhaustive negatives and exhaustive annotations.
Despite strong performance on classes with many and medium, the upper bound for classes with few examples remains low (4.21 AP and 3.93 AP). We posit that it is very hard to capture the correct semantics of nuImages’ rare categories only using $K$-shots. We observe similar trends for the 5 and 30-shot cases and present further analysis in the supplement. Given the success of training with pseudo-negatives, a natural next-step is to train with pseudo-positives. Our preliminary results suggest that incorporating pseudo-positives does not provide significant improvement over simply training with pseudo-negatives. We posit that training with incorrect pseudo-positives may incur a higher penalty than training with incorrect pseudo-negatives. This is a promising direction for future work.
Approach 10 Shots: Average Precision (AP)
@@ 4.4 Oracle Performance Analysis
Detic (Zero-Shot) [49] 14.26 27.28 16.88 2.36
   + Fine-Tuning 15.53 26.01 19.93 3.88
   w/ FedLoss 15.57 27.20 20.09 2.89
   w/ Inverse FedLoss 15.89 27.56 20.19 3.42
   w/ Pseudo-Negatives 16.67 29.15 20.66 3.90
   w/ True Negatives (Oracle) 16.99 29.60 20.93 4.21
   w/ Exhaustive Annotations (Oracle) 18.51 33.51 22.35 3.93
4.5 Fine-Tuning Ablation
We explore different fine-tuning strategies for training Detic with few-shot annotations. We broadly divide Detic’s architecture into four components: Backbone, Region Proposal Network (RPN), Box Regressor, and Classifier. We ablate the impact of freezing different components and present results in Table 5.
Intuitively, as we have limited training data, we attempt to fine-tune a minimal number of parameters. Initializing and freezing the classifier head with CLIP embeddings corresponding to class names provides the most significant improvement. Prior works that fine-tune vision-only models have no notion of language embeddings and therefore must train classifiers from scratch. In contrast, Detic can represent class names using CLIP embeddings and can more easily adapt to novel categories with few examples. We find that freezing the backbone, RPN and classifier head with CLIP embeddings, and training the classifier projection layer and box regressor performs the best. Importantly, as Detic has been trained on large-scale datasets, its RPN can easily localize novel objects without fine-tuning.

### v3

@@ 4 Experiments
Table 1: VLM Zero-Shot Inference Is a Strong FSOD Baseline. Zero-shot inference with VLMs like GroundingDINO resoundingly outperforms state-of-the-art FSOD methods on the COCO FSOD benchmark, motivating the need to re-frame FSOD to embrace foundation models.
@@
AP Base AP Novel AP
FRCN-ft-full [yan2019meta] 18.6 20.6 12.5
FRCN-BCE [yan2019meta] 30.2 36.8 10.3
TFA w/ fc [wang2020frustratingly] 29.3 34.5 13.5
TFA w/cos [wang2020frustratingly] 29.9 35.3 13.6
MPSR [wu2020multi] 17.1 18.1 14.1
Meta-RCNN [yan2019meta] 7.8 7.1 9.1
FsDetView [xiao2022few] 10.0 9.3 12.0
Retentive R-CNN [fan2021generalized] 32.9 39.3 13.8
DiGeo [ma2023digeo] 33.1 39.4 14.2
GroundingDINO (Zero-Shot) [liu2023grounding] 48.3 46.3 54.3
@@ 4.1 Zero-Shot Inference Is A Strong FSOD Baseline
We compare state-of-the-art FSOD methods with zero-shot inference from GroundingDINO [liu2023grounding] on COCO in Table 1. Surprisingly, GroundingDINO outperforms DiGeo [ma2023digeo] by 16.2% AP averaged across both base and novel categories despite never being trained on COCO images. GroundingDINO’s impressive performance is due to its large-scale multi-modal pre-training on Objects365 [shao2019objects365], GoldG [goldg] and Cap4M [li2021grounded]. It is worth noting that GroundingDINO achieves higher AP on novel classes than base, suggesting that novel classes in existing benchmarks (e.g. car and person) are actually not rare in the real world. Therefore, FSOD benchmarks should be re-framed to reflect real-world applications, motivating our setup.
@@
Table 2: Impact of Large-Scale Pre-Training and Language. We repurpose nuImages for FSOD following the dataset creation process established by [wang2020frustratingly]. We group categories by frequency into many, medium and few examples per class [peri2023towards, ma2023longtailed]. We fine-tune TFA on $K$ examples, but find low performance, $<3$AP. However, by simply pre-training on more data (LVIS, COCO and ImageNet-21K) and leveraging language cues via a CLIP classifier, 5-shot performance improves from 2.02 AP to 15.12 AP. However, rare (or few) classes like strollers, pushable-pullable, and debris remain challenging, motivating our task of VLM alignment.
@@
   TFA [wang2020frustratingly] w/ COCO-base 1.33 2.78 1.43 0.23
   TFA [wang2020frustratingly] w/ LVIS-base 2.02 1.69 4.08 0.58
   TFA [wang2020frustratingly] w/ LVIS,IN-21K, 15.12 22.74 18.99 4.25
@@
   TFA [wang2020frustratingly] w/ COCO-base 1.21 2.55 1.19 0.31
   TFA [wang2020frustratingly] w/ LVIS-base 2.27 2.05 4.51 0.58
   TFA [wang2020frustratingly] w/ LVIS,IN-21K, 16.09 25.46 20.00 3.73
@@
   TFA [wang2020frustratingly] w/ COCO-base 1.14 2.81 0.84 0.23
   TFA [wang2020frustratingly] w/ LVIS-base 2.23 1.48 4.98 0.45
   TFA [wang2020frustratingly] w/ LVIS,IN-21K, 17.22 25.98 21.64 4.78
@@ 4.2 Foundational FSOD with nuImages
In the context of foundational models, we argue that partitioning datasets into base and novel classes no longer makes sense. Instead, FSOD methods should only train on $K$-shot annotations for $C$ target classes, and also evaluate performance on these $C$ classes. We pre-train TFA [wang2020frustratingly] on diverse datasets and fine-tune on $K$ examples per class and highlight model performance in Table 2. We train two variants of TFA trained on COCO-base and LVIS-base and fine-tune both models on $K$ examples of the nuImages classes. Surprisingly, both variants of TFA achieve less than $3$ AP (cf. Table 2). We posit that this is largely due to poor classification performance. Since both LVIS and COCO classes do not significantly overlap with nuImages classes, learning a classifier from few examples is extremely difficult. However, we find that simply re-training TFA with a frozen CLIP-based classifier (similar to Detic) dramatically increases performance, reiterating the utility of language and web-scale pre-training in data-constrained settings.
Table 3: Empirical Analysis of Baselines (10-Shots) on our Benchmark. We evaluate popular VLMs on the nuImages FSOD Benchmark and find that MQ-GLIP performs the best among all baseline models. Notably, it achieves 17.0 mAP zero-shot language-only performance, and achieves 21.4 mAP via zero-shot multi-modal prompting averaged over all classes. Remarkably, our 2024 competition winners further improved performance to 45.4 mAP, beating our best baseline by 24.0%.
Approach Backbone Pre-Train Data Average Precision (AP)
All Many Med Few
Zero-Shot Detection
   RegionCLIP [zhong2022regionclip] RN50 CC3M 2.50 3.20 3.80 0.40
   Detic [zhou2022detecting] SWIN-B LVIS, COCO, IN-21K 14.40 25.83 16.59 2.32
   GroundingDINO [liu2023grounding] SWIN-T Objects365, GoldG, Cap4M 12.05 17.29 15.45 3.72
   GLIP [li2021grounded] SWIN-L FourODs,GoldG,Cap24M 17.01 23.36 19.86 8.40
   MQ-GLIP-Text [xu2024multi] SWIN-L Objects365, FourODs, GoldG, Cap24M 17.01 23.36 19.85 8.41
Prompt Engineering
   Detic [zhou2022detecting] SWIN-B LVIS, COCO, IN-21K 14.92 26.48 17.29 2.53
   GLIP [li2021grounded] SWIN-L FourODs, GoldG, Cap24M 17.15 23.82 19.36 9.02
Standard Fine-Tuning
   RegionCLIP [zhong2022regionclip] RN50 CC3M 3.86 6.08 5.13 0.54
   Detic [zhou2022detecting] SWIN-B LVIS, COCO, IN-21K 16.09 25.46 20 3.73
Federated Fine-Tuning (Ours)
   Detic [zhou2022detecting] SWIN-B LVIS, COCO, IN-21K 17.24 28.07 20.71 4.18
   Detic [zhou2022detecting] w/ Prompt Engineering SWIN-B LVIS, COCO, IN-21K 17.71 28.46 21.14 4.75
Language Prompt Tuning
   GLIP [li2021grounded] SWIN-L FourODs,GoldG,Cap24M 19.41 22.18 25.16 10.39
Visual Prompting
   MQ-GLIP-Image [xu2024multi] SWIN-L Objects365,FourODs,GoldG,Cap24M 14.07 24.39 15.89 3.34
Multi-Modal Prompting
   MQ-GLIP [xu2024multi] SWIN-L Objects365,FourODs,GoldG,Cap24M 21.42 32.19 23.29 10.26
Multi-Modal Chat Assistants
   GPT-4o Zero-Shot Classification [achiam2023gpt] Private Private 9.95 16.81 12.11 1.71
CVPR 2024 Competition Results
   PHP_hhh Private Private 45.35 64.25 53.43 20.19
   NJUST KMG SWIN-L Objects365V2, OpenImageV6, GoldG, V3Det, COCO2014, COCO2017, LVISV1, GRIT, RefCOCO, RefCOCO+, RefCOCOg, gRef-COCO 32.56 50.21 34.87 15.16
   zjyd_cxy_vision SWIN-L Objects365V2, COCO2017, LVIS, GoldG, VG, OpenImagesV6, V3Det, PhraseCut, RefCOCO, RefCOCO+, RefCOCOg, gRef-COCO 31.57 46.59 33.32 17.03
4.3 Empirical Analysis of Results
We evaluate several popular VLMs on the nuImages Foundational FSOD (10-shots) benchmark and present salient insights from Table 3 below.
Zero-Shot Detection. Somewhat unsurprisingly, we find that (i) greater pre-training data scale and diversity, along with (ii) larger backbones result in better zero-shot performance. Notably, GLIP achieves 17.01% zero-shot performance, surpassing all other methods trained with less data and smaller backbones.
Prompt Engineering. We attempt to improve zero-shot performance using synonyms for class names derived from the annotator textual instructions. We see minor improvements (e.g. Detic improves from $14.40$ mAP $\rightarrow 14.92$ mAP), indicating that leveraging rich textual descriptions beyond class names can improve concept alignment.
Federated Fine-Tuning. Standard fine-tuning is sub-optimal for FSOD, as all unannotated classes are treated as negatives. Therefore we use our zero-shot predictions to generate pseudo-labels on training images. We extract pseudo-negatives based on these pseudo-labels by identifying classes not in each image (by using detector confidence scores), and leverage pseudo-negatives in our fine-tuning. Notably, we improve over Detic’s standard fine-tuning by $1.15$ mAP ($16.09$ mAP $\rightarrow 17.24$ mAP).
Multi-Modal Prompting. We observe that Multi-Modal Prompting (MQ-GLIP) achieves the best performance ($21.42$ mAP) out of all open-source methods in Table 3. We attribute this to its large pre-trained dataset, bigger backbone (SWIN-L) and multi-modal prompts used during inference. Notably, the benefit of multi-modal prompts can be seen by comparing MQ-GLIP ($21.42$ mAP) against MQ-GLIP-Image ($14.07$ mAP), which uses visual prompting and MQ-GLIP-Text ($17.01$ mAP), which uses language prompting. Interestingly, MQ-GLIP does not require gradient-based fine-tuning, which differs from all existing conventional few-shot methods. Therefore, we posit that future few-shot methods should further explore in-context learning. Just as multi-modal annotator instructions aid human annotator alignment, we find that multi-modal prompting significantly improves VLM concept alignment.
Multi-Modal Chat Agents. Given the strong performance of GPT-4o for general visual understanding, we repurpose it for our task by prompting the model to re-classify image patches from Detic’s RPN. Specifically, we ask GPT-4o to predict a class and confidence for each image crop. Surprisingly, we observe reasonable performance (9.95 mAP) despite GPT-4o not being trained as an object detector, emphasizing the importance of the scale of pre-training data. We explore iterative prediction refinement in the supplement.
CVPR 2024 Challenge. Finally, we highlight our top three submissions (out of six participants) from the inaugural Foundational FSOD challenge. Notably all top performers beat our baselines, with the winning team achieving $45.35$ AP! We discuss more details in the supplement.
Figure 3: Visualizing Random and Best Split. In the top row, we visualize the 5-shot training examples of strollers from a random split. Similarly, we visualize the 5-shot training examples from the best split in the bottom row. We observe that strollers in the random split are often occluded, small in size and blurry, making few-shot learning harder. On the other hand, the best split examples are larger, have better visual quality and are relatively un-occluded. This visual difference directly translates into better few-shot performance. We achieve $\mathbf{13.09}$ Stroller AP for the random split and $\mathbf{18.54}$ Stroller AP for the best split. We show a more comprehensive evaluation in Table 4.
Table 4: Random Split vs “Best” Split. We construct the “best" split by selecting per-class few-shot examples that lead to the highest performance on a held-out set. Unsurprisingly, the best split performs better than any random split, especially for very limited data settings (e.g. $5$-shot detection). This evaluation setting closely mimics how human annotators are “aligned” to target concepts, since annotator guides are constructed using hand-picked iconic visual examples.
Approach Average Precision (AP)
@@ 4.3 Empirical Analysis of Results
Detic (Zero-Shot) [zhou2022detecting] 14.40 25.83 16.59 2.32
Detic w/ Federated Fine-Tuning (5-shots, Random Split) 16.58 27.12 19.71 4.13
Detic w/ Federated Fine-Tuning (5-shots, Best Split) 18.30 28.66 21.81 5.56
Detic w/ Federated Fine-Tuning (10-shots, Random Split) 17.24 28.07 20.71 4.18
Detic w/ Federated Fine-Tuning (10-shots, Best Split) 18.24 28.63 22.00 5.19
Detic w/ Federated Fine-Tuning (30-shots, Random Split) 18.64 29.13 22.44 5.46
Detic w/ Federated Fine-Tuning (30-shots, Best Split) 18.75 28.07 23.18 5.81
4.4 Analysis of Iconic Few-Shot Images
The specific examples used during few-shot fine-tuning significantly impacts target class performance [wang2020frustratingly]. However, prior work constructs few-shot splits by randomly sampling $K$ examples per class. In contrast, when creating annotator instructions, selecting the right examples to “align" human annotators [chang2023thinking] to subtle aspects of the target concept is carefully considered. To more closely match VLM concept alignment with human annotator alignment, we design a simple algorithm to construct the best $K$-shot split for fine-tuning. This allows us to understand which examples are most informative and measure an upper bound in performance.
We construct our best split by picking examples corresponding to the best class-wise performance, based on the evaluation of each split on a held-out validation set. For instance, out of $3$ random splits for the $5$-shot task, one may pick car examples from split $1$, bicycle from split $3$ and debris from split $2$. In Table 4, we observe that the best-split performance is always better than its random counterpart. As expected, the choice of examples in $5$-shot case is more important than the $30$-shot case ($1.72$ AP difference for $5$-shot vs $0.11$ AP for $30$-shots). We visualize the difference in the splits for strollers in nuImages (cf. Figure 3). Unsurprisingly, iconic examples are large and unoccluded.
4.5 Limitations and Future Work
Despite using VLMs pre-trained on large-scale datasets, we find that performance for rare categories (defined by the cardinality of each class in the original dataset) is considerably lower than for common classes. We posit that VLMs are pre-trained with imbalanced data which includes many examples of common categories like truck but few examples of rare categories like stroller. Our work does not explicitly improve detection performance on rare classes. Interestingly, since VLMs like Detic [zhou2022detecting], GLIP [li2021grounded], and GroundingDINO [liu2023grounding] are trained with different data sources, each model has dramatically different zero-shot performance on novel categories like stroller. Ensembling predictions from different VLMs may yield better detection accuracy for rare categories. In addition, although our work motivates the use of rich textual descriptions found in instructions for multi-modal alignment, our current results use only nouns (class names and synonyms) as text prompts.
Benchmarking in the Era of Foundation Models. Although we argue that pre-training on large-scale data will be the key enabler for generalization to the open-world, understanding how to appropriately benchmark such methods remains challenging. It is readily accepted that in order to accurately evaluate generalization, one should not train on test data. However, it is difficult to guarantee that foundation models have never seen our specific test data. We address this in our challenge by explicitly prohibiting participants from training on nuImages (and nuScenes). However, should we allow participants to train on similar in-domain data (e.g., other autonomous vehicle datasets such as Argoverse [Argoverse2])? We argue ‘yes’! With enough scale, novel test examples may still be similar to the training set.
Out-of-Domain Benchmarks. Another way to address benchmarking is to collect test scenarios that are designed to be dissimilar from internet images. For example, out-of-domain images may include medical data (though foundational performance is still surprisingly effective [Wang2022MedCLIPCL]). We admittedly did not take this route, since urban imagery is similar to images found online and arguably many applications of interest fall under this category. Moreover, there exist ample opportunity for technical innovation in this setting (as suggested by our CVPR 2024 challenge results!). Alternatively, one can manually collect and sequester images that will never be released on the internet. Since ensuring privacy may itself be challenging, yet another approach is to leverage the continual learning paradigm, where new test sets are continually constructed over time.
Comparing Models. Fairly comparing foundation models requires careful consideration. Although accuracy is a valuable metric, it is intrinsically tied to the scale of pre-training data and model architecture. Notably, the LLM community already ranks models via a Pareto frontier of accuracy vs. parameter count. We advocate for a similar approach for Foundational FSOD that considers backbone architecture (e.g. ResNet-50 vs. Swin-B) and pre-training datasets (e.g. CC4M, GoldG, LVIS).
5 Conclusion
We revisit few-shot object detection (FSOD) with vision-language models (VLMs) and find that zero-shot inference from web-scale VLMs significantly outperforms leading FSOD methods. However, such foundational models do not fully address few shot recognition because of the concept alignment problem; particular concepts in target applications may be different than their use on web-scale datasets. Just as human annotators require concept alignment via multi-modal text and visual examples, we argue that VLMs should be aligned with such few-shot data, formalizing the problem of Foundational FSOD.